The electric field at a point in space is E⃗  =( 400i^+ 700 j^)N/C.

What is the x-component of the electric force on a proton at this point?

What is the y-component of the electric force on a proton at this point?

What is the x-component of the electric force on an electron at this point?

What is the y-component of the electric force on a electron at this point?

What is the magnitude of the proton's acceleration?

What is the magnitude of the electron's acceleration?
Answer #1

a) Fx = 1.6 * 10-19 * 400 = 6.4 * 10-17 N

b) Fy = 1.6 * 10-19 * 700 = 1.12 * 10-16 N

c) Fx = -1.6 * 10-19 * 400 = -6.4 * 10-17 N

d) Fy = -1.6 * 10-19 * 700 = -1.12 * 10-16 N

e) Fnet = sqrt [(6.4 * 10-17)2 + (1.12 * 10-16)2] = 1.28 * 10-16 N

ap = Fnet / mp = (1.28 * 10-16 ) / (1.67 * 10-27) = 7.7 * 1010 m/s2

f) ae = Fnet / me = (1.28 * 10-16 ) / (9.11 * 10-31) = 1.4 * 1014 m/s2